OpenStudy (anonymous):

How would you solve: .5/sin(30)=c/sin(90) @Mertsj @tomo

OpenStudy (mertsj):

Do you know the sin of 30?

OpenStudy (anonymous):

.5 @Mertsj

OpenStudy (anonymous):

.5/sin(30)=c/sin(90) multiply both sides by 2 to get rid of the .5 1/sin(30) = 2c/sin(90) what is sin(90) you can put it in a calculator and get 1 so you have 1/sin(30) = 2c sin(30) = 1/2 so you have 1/(1/2) = 2c so 2 = 2c c = ?

OpenStudy (anonymous):

So c would be equal to 1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay and one more :) .5/sin(30)=a/sin(60) would you do the same steps?

OpenStudy (anonymous):

yes, just solve for a.

OpenStudy (anonymous):

awesome! thanks :3